# Herding Unmasked: Insights into Cryptocurrencies, Stocks and US ETFs

### v1

@@ 4.1 Herding Detection on Each Type of Investment Vehicle
In summary, we found that investors in all three types of investment vehicle showed evidence of herding during the Covid-19 pandemic sub-period. This is reasonable since the serious pandemic directly affected the global economy and all financial markets. However, other sub-periods reveal different herding patterns between the cryptocurrency market and traditional markets. In particular, herding behavior was found only in stocks and US ETFs throughout 2019, while this phenomenon disappeared in the cryptocurrency market. Indeed, herding during this time appeared to be caused by the trade war between China and the US which was relevant to macroeconomic factors and directly influenced various companies in different sectors. Thus, traditional markets were undoubtedly affected the most. On the other hand, only the cryptocurrency market displayed a herding signal during the Bull market time. This can be explained by the fact that cryptocurrency investors are often swayed more by positive news and positive market returns than negative ones [88]. Whereas, traditional markets are typically more responsive to distressed market conditions than booms [117, 19]. In particular, when traditional markets experience an upward trend, professional investors are found to account for the majority portion [4]. Thus, they tend to make investment decisions by themselves which prevents them from performing as a herd. For the cryptocurrency market, however, its investors are more likely to be naïve and make irrational decisions, based on word of mouth as well as existing news on social media [57, 8]. Thus, positive news surrounding cryptocurrencies can lure inexperienced investors with the promise of substantial profits.
@@ 4.2 Herding Detection on a Combination of Cryptocurrencies, Stocks and US ETFs
Furthermore, we provide information about sector distribution for each community. Specifically, each asset in a community is categorized into either cryptocurrency, or US ETF, or one of the 11 stock sectors discussed in section 3. Subsequently, the proportion of each sector is calculated in each community. Finally, for each community in each sub-period, we can find whether herding behavior exists in those assets and its corresponding sector distribution.
@@ 4.2.1 Community Structures for a Combination of Cryptocurrencies, Stocks and US ETFs
We first introduce the community structure obtained from a combination of cryptocurrencies, stocks and US ETFs in each sub-period, which is displayed in Figs 4-5. Each node in the graph represents an asset, while an edge between two nodes represents their correlation. If there is no edge between two nodes, then their correlation is negligible compared to other surrounding linked nodes. Each color represents a community in which assets with the same color belong to one community, meaning that they behave similarly and their intercorrelations are stronger than those of correlations with other nodes in different communities. The longer the path between two nodes, the more dissimilar they are.
(a) Pre-Covid-19 (04/2019 - 12/2019)
(b) Covid-19 pandemic (01/2020 - 06/2020)
(c) Post-Covid-19 (07/2020 - 12/2020)
Figure 4: Time-varying community structures of cryptocurrencies, stocks and US ETFs during: a) Pre-Covid-19, b) Covid-19 pandemic, c) Post-Covid-19. More information in each sub-figure can be seen at Clarification 3..
(a) Bull market time (01/2021 - 04/2022)
(b) Ukraine-Russia Conflict (05/2022 - 05/2023)
Figure 5: Time-varying community structures of cryptocurrencies, stocks and US ETFs during: a) Bull market period, b) Ukraine-Russia conflict. More information in each sub-figure can be seen at Clarification 3..
As can be seen in Figs 4-5, the community structures change over time, indicating the variation in correlations among assets over time. Notably, we found some common patterns out of these five different structures. Firstly, cryptocurrencies form several separate communities but these are next to each other, meaning that there is a strong correlation between cryptocurrencies themselves. Furthermore, there is still a distinction between the cryptocurrency market and traditional markets since most cryptocurrencies do not form communities with stocks and US ETFs, an example of this characteristic is highlighted within a purple-shaped oval in each figure. Secondly, it can be seen that stocks and US ETFs tend to belong to the same community. Interestingly, US ETFs seem to be the centre of a community, linked by stock assets, indicating that they are the most influential nodes in a community, an example of this characteristic is highlighted within a black-shaped oval in each figure. Thirdly, there are some stock sectors in which stocks within the same sector and ETFs that contain these stocks tend to be in the same communities, namely Energy, Healthcare, Information Technology and Finance. On the other hand, this feature does not appear in the remaining sectors. Fourthly, we also notice that S&P-related ETFs tend to belong to one community throughout the considered sub-periods.
@@ 4.2.2 Herding Behavior at Community Level
Figs 6-10 illustrate the herding behavior results and sector distribution of each community in each sub-period. The assets in each community are listed in Appendix E. At first glance, it is clear that the first two sub-periods look rather similar in terms of the density of herding signals, as most communities reveal herding behavior in Figs 6-7. This is contrary to the last three sub-periods (Figs 8-10) since these predominantly display a dearth of herding. Perhaps, the high density of herding during the pre-and the peak of the pandemic was caused by two major events, as mentioned previously, the US-China trade war throughout 2019 and the contagious disease Covid-19 which started early in 2020, respectively. Whereas, once the pandemic started to be controlled and the global economy recovered from the loss, financial markets resumed their normal behavior. Thus, investors stopped mimicking each other and began to make their own objective decisions, and the herding was lessened accordingly.
(a) Herding Detection Results
(b) Sector Distribution
@@
Figure 6: Herding detection results (a) and sector distribution (b) of each community during the Pre-Covid-19 period. More information on each sub-figure can be seen at Clarification 2..
(a) Herding Detection Results
(b) Sector Distribution
@@
Figure 7: Herding detection results (a) and sector distribution (b) of each community during the Covid-19 pandemic period. More information on each sub-figure can be seen at Clarification 2..
(a) Herding Detection Results
(b) Sector Distribution
Figure 8: Herding detection results (a) and sector distribution (b) of each community during the Post-Covid-19 period. More information on each sub-figure can be seen at Clarification 2..
(a) Herding Detection Results
(b) Sector Distribution
@@
Figure 9: Herding detection results (a) and sector distribution (b) of each community during the Bull market period. More information on each sub-figure can be seen at Clarification 2..
(a) Herding Detection Results
(b) Sector Distribution
@@
Figure 10: Herding detection results (a) and sector distribution (b) of each community during the Ukraine-Russia conflict period. More information on each sub-figure can be seen at Clarification 2..
@@
We notice from Figs 6-10 that cryptocurrencies are still mostly isolated from traditional markets to some extent since they are not distributed into multiple communities. Instead, they are close to each other and tend to form 2 to 3 large communities among themselves. Observing these communities throughout different sub-periods, it is clear that they are herded most of the time. Compared to the herding detection results in the cryptocurrency market as discussed in the previous section, it is clear that without critical events impacting the entire cryptocurrency market, herding still exists in most cryptocurrencies, but at a community level, i.e herding behaviour is pronounced in different subsets of cryptocurrencies, while there was an absence of herding signals when looking at all cryptocurrencies at once, as shown in the previous section. It is challenging to find the reasons causing the herding in such cases since there was no major event that could influence the market [7, 89, 119]. However, this market seems to be mainly driven by investors’ emotions [94] and accessible news on online platforms [21]. In addition, cryptocurrency investors tend to adopt a short-term investment strategy and depend on price movements of the coins rather than reliable factors that are seen in traditional markets [90]. Consequently, the herding in those cryptocurrency communities could arise from either the potential profits or the fear of losing money at that time, encouraging investors to buy or sell. Moreover, investors might be aware of the co-movements between different cryptocurrencies. Thus, when the opportunity to gain profits (or the selling time) arrives, they tend to buy (or sell) the coins that have the same price movements to get higher profits (or mitigate the loss). This could explain why cryptocurrencies formed 2 separate communities in the Pre-Covid-19 and also the Bull market time, with both communities revealing herding behavior, while there was no herding when we detected the whole cryptocurrency market in the previous section.
@@
Regarding other communities, each of them has a mixture of US ETFs and stocks where the US ETFs tend to act as a basket of most stocks in that community. The herding in these communities follows the same rule as the cryptocurrency communities since each sub-period reveals significant herding behavior in several communities. Remarkably, we obtained several valuable findings from the herding results of these stocks and US ETFs. In particular, there are 4 stock sectors (Energy, Healthcare, IT and Finance) in which for each sector, its assets and relevant ETFs tend to form communities among themselves, throughout the examined period. Moreover, the herding signals found in each of these sectors coincide with the timing of economic and political events that specifically influence that sector.
@@
  • Technology Sector: Technology stocks are grouped into several communities and reveal herding behavior in three sub-periods, including Pre-Covid-19 (coinciding with the US-China trade war [124, 125, 126, 126]), Covid-19 pandemic (coinciding with the downturn time caused by the pandemic followed by a global economic recession) and the Ukraine-Russia conflict (coinciding with the time when Technology companies were laying off workers [127, 128, 129] followed by the AI boom that gave an impetus to the Technology sector [130, 131]).
@@
  • Healthcare Sector: Herding behavior was found in most healthcare stocks and ETFs during three sub-periods: the Covid-19 pandemic, the Post-Covid-19 and the Bull market. This herding behavior likely stems from the great attraction of investors toward pharmaceutical companies due to their positive performance during the pandemic. In particular, dozens of pharmaceutical companies began vaccine research from early 2020 due to the high demand for vaccinations for the entire world [132, 133, 134]. Moreover, other pharmaceutical companies, though not participating in producing vaccines, produced critical elements that were necessary for vaccine production such as nanoparticles or medical devices for hospitals and daily usage such as the Covid test kits. These contributions also pushed their performance during this time [134, 135].
@@
  • Financial Sector: In this study, the financial sector mainly consists of banking firms. We found that this industry was exposed to herding behavior in the first four sub-periods, except for the Ukraine-Russia conflict time. The financial sector is strongly related to economic conditions and reflects the performance of the economy, especially the banking firms since they have the responsibility of preserving the economic stability of a country and also keeping the country away from financial crises through various schemes like reducing interest rates [136], imposing new monetary policy comes from FED [137] and cutting companies’ liquidity risks [138]. Consequently, economic uncertainty forces the financial sector to experience distress. Indeed, the global economy had been experiencing ups and downs continually through out the herding periods in this stock sector.
For the remaining sectors (i.e. Industry, Consumer Discretionary and Consumer Staples), we found that assets within the same sector never group into a large community in all sub-periods, even during the Covid-19 pandemic. Instead, they are broken down and distributed almost equally into several communities. This result indicates that investors did not perform herding on a sector basis in these sectors, even during major events. Remarkably, however, we do notice that there is simultaneous herding behavior between the Technology sector and Standard &\&& Poor (S&\&&P)-related ETFs, as shown in Fig 6 (Communities 6 and 7 for Technology stocks, Community 10 for S&\&&P-related ETFs), Fig 7 (Community 5 for Technology stocks, Community 12 for S&\&&P-related ETFs), and Fig 10 (Community 5 for Technology stocks, Community 8 for S&\&&P-related ETFs). Specifically, when the herding behavior is pronounced in the Technology sector, a set of ETFs tracking the S&P-related stock indices also shows this phenomenon, including IJR, IJS, IJT, IWM, SLYS, SLYV and SLY. These are ETFs representing the aggregate performance of small-cap stocks published by Standard & Poor. This finding is surprising since all the technology stocks used in this work are large-cap ones but their behavior is unexpectedly associated with ETFs tracking the performance of small-cap stocks. This relationship between the technology sector and these ETFs will be considered in future work.
@@
Furthermore, we acknowledge that some communities are mixed by multiple sectors and reveal herding, but it is unclear why the herding behavior exists in those cases as there seems not to be a rule/convention for these cases. However, most of such communities appear during the economic uncertainty times, i.e. during the trade war (Fig 6) and the Covid-19 pandemic (Fig 7). By contrast, this phenomenon was not pronounced from the Bull market time onwards (Figs 9-10). Thus, it seems to be more likely that negative market conditions resulted in this phenomenon. However, since we don’t have a confident justification for this phenomenon, we will leave this result for future work.
@@
From this experimental result, we can conclude that herding behavior should be observed from two perspectives, one from the entire market and the other from separate subsets (communities) within the market. This is because different perspectives provide us with different aspects of the market. Specifically, herding behavior occurs widely in most assets of a particular type of investment vehicle when there is a critical event that directly impacts the market. Hence, most assets involved are influenced and thus investors who are trading those assets tend to perform herding as a reaction to the event. This is why we found the herding in traditional markets during the US-China trade war, the herding in all markets during the peak of the Covid-19 pandemic and the herding in the cryptocurrency market during the Bull market time, while the remaining sub-periods show non-herding results. By contrast, we found that herding behavior occurs constantly in all types of investment vehicle but at a smaller scale, i.e. community level, as shown in Figs 6-10. In addition, herding behavior found in these communities mainly comes from specific events that solely impact that community. Subsequently, although herding might not exist in a whole market, it can be found by looking at a particular community/sector.

### v2

@@ 4.1 Herding Detection on Each Type of Investment Vehicle
In summary, we found that investors in all three types of investment vehicle showed evidence of herding during the Covid-19 Pandemic sub-period. This is reasonable since the serious pandemic directly affected the global economy and all financial markets. However, most remaining sub-periods reveal different herding patterns between the cryptocurrency market and traditional markets. In particular, herding behavior was found only in stocks and US ETFs throughout 2019, while this phenomenon disappeared in the cryptocurrency market. Indeed, herding during this time appeared to be caused by the trade war between China and the US which was relevant to macroeconomic factors and directly influenced various companies in different sectors. Thus, traditional markets were undoubtedly affected the most. On the other hand, only the cryptocurrency market displayed a herding signal during the first Bull Time and second Ukraine-Russia Conflict sub-periods. This can be explained by the fact that cryptocurrency investors are often swayed more by positive news and positive market returns than negative ones [90]. On the other hand, traditional markets are typically more responsive to distressed market conditions than booms [131, 19]. In particular, when traditional markets experience an upward trend, professional investors are found to account for the majority portion [4]. Thus, they tend to make investment decisions by themselves which prevents them from performing as a herd. For the cryptocurrency market, however, its investors are more likely to be naïve and make irrational decisions, based on word of mouth as well as existing news on social media [58, 8]. Thus, positive news surrounding cryptocurrencies can lure inexperienced investors with the promise of substantial profits.
@@ 4.2 Herding Detection on a Combination of Cryptocurrencies, Stocks and US ETFs
Furthermore, we provide information about sector distribution for each community. Specifically, each asset in a community is categorized into either cryptocurrency, or US ETF, or one of the 11 stock sectors discussed in section Methodology and Data Collection. Subsequently, the proportion of each sector is calculated in each community. Finally, for each community in each sub-period, we can find whether herding behavior exists in those assets and its corresponding sector distribution.
@@ 4.2.1 Community Structures for a Combination of Cryptocurrencies, Stocks and US ETFs
We first introduce the community structure obtained from a combination of cryptocurrencies, stocks and US ETFs in each sub-period, which is displayed in Figs 2-8. Each node in the graph represents an asset, while an edge between two nodes represents their correlation. If there is no edge between two nodes, then their correlation is negligible compared to other surrounding linked nodes. Each color represents a community in which assets with the same color belong to one community, meaning that they behave similarly and their intercorrelations are stronger than those of correlations with other nodes in different communities. The longer the path between two nodes, the more dissimilar they are.
As can be seen in Figs 2-8, the community structures change over time, indicating the variation in correlations among assets over time. Notably, we found some common patterns out of these seven different structures. Firstly, cryptocurrencies form several separate communities but these are next to each other, meaning that there is a strong correlation between cryptocurrencies themselves. Furthermore, there is still a distinction between the cryptocurrency market and traditional markets since most cryptocurrencies do not form communities with stocks and US ETFs, an example of this characteristic is highlighted within a purple-shaped oval in each figure. Secondly, it can be seen that stocks and US ETFs tend to belong to the same community. Interestingly, US ETFs seem to be the centre of a community, linked by stock assets, indicating that they are the most influential nodes in a community, an example of this characteristic is highlighted within a black-shaped oval in each figure. Thirdly, there are some stock sectors in which stocks within the same sector and ETFs that contain these stocks tend to be in the same communities, namely Energy, Healthcare, Information Technology and Financials. On the other hand, this feature does not appear in the remaining sectors. Fourthly, we also notice that S&P-related ETFs tend to belong to one community throughout the considered sub-periods.
Figure 2: Community structure of cryptocurrencies, stocks and US ETFs during the Pre-Covid-19 period. More information can be seen at S2 Text.
Figure 3: Community structure of cryptocurrencies, stocks and US ETFs during the Covid-19 Pandemic period. More information can be seen at S2 Text.
Figure 4: Community structure of cryptocurrencies, stocks and US ETFs during the first Bull Time period. More information can be seen at S2 Text.
Figure 5: Community structure of cryptocurrencies, stocks and US ETFs during the second Bull Time period. More information can be seen at S2 Text.
Figure 6: Community structure of cryptocurrencies, stocks and US ETFs during the third Bull Time period. More information can be seen at S2 Text.
Figure 7: Community structure of cryptocurrencies, stocks and US ETFs during the first Ukraine-Russia period. More information can be seen at S2 Text.
Figure 8: Community structure of cryptocurrencies, stocks and US ETFs during the second Ukraine-Russia period. More information can be seen at S2 Text.
@@ 4.2.2 Herding Behavior at Community Level
Figs 9-15 illustrate the herding behavior results (Figs 9a-15a) and sector distribution (Figs 9b-15b) of each community in each sub-period. The assets in each community are listed in S4 Table.. At first glance, it is clear that the first two sub-periods look rather similar in terms of the density of herding signals, as most communities reveal herding behavior in Figs 9a and 10a. This is contrary to the last five sub-periods (Figs 11a-15a) since these predominantly display a dearth of herding. Perhaps, the high density of herding during the pre-and the peak of the pandemic was caused by two major events, as mentioned previously, the US-China trade war throughout 2019 and the contagious disease Covid-19 which started early in 2020, respectively. Whereas, once the pandemic started to be controlled and the global economy recovered from the loss, financial markets resumed their normal behavior. Thus, investors stopped mimicking each other and began to make their own objective decisions, and the herding was lessened accordingly.
@@
Figure 9: Herding detection results (a) and sector distribution (b) of each community during the Pre-Covid-19 period. More information on each sub-figure can be seen at S1 Text..
@@
Figure 10: Herding detection results (a) and sector distribution (b) of each community during the Covid-19 Pandemic period. More information on each sub-figure can be seen at S1 Text..
@@
Figure 11: Herding detection results (a) and sector distribution (b) of each community during the first Bull Time period. More information on each sub-figure can be seen at S1 Text..
@@
Figure 12: Herding detection results (a) and sector distribution (b) of each community during the second Bull Time period. More information on each sub-figure can be seen at S1 Text..
Figure 13: Herding detection results (a) and sector distribution (b) of each community during the third Bull Time period. More information on each sub-figure can be seen at S1 Text..
Figure 14: Herding detection results (a) and sector distribution (b) of each community during the first Ukraine-Russia Conflict period. More information on each sub-figure can be seen at S1 Text..
Figure 15: Herding detection results (a) and sector distribution (b) of each community during the second Ukraine-Russia Conflict period. More information on each sub-figure can be seen at S1 Text..
@@
We notice from Figs 9b-15b that cryptocurrencies are close to each other and tend to form 2 to 3 large communities among themselves. Observing these communities throughout different sub-periods, there is evidence of herding most of the time, as shown in Figs 9a-15a. Compared to the herding detection results in the cryptocurrency market displayed in Tables 3 and 4, it is clear that without critical events impacting the entire cryptocurrency market, herding still exists in most cryptocurrencies, but at a community level, i.e herding behaviour is pronounced in different subsets of cryptocurrencies, while there was an absence of herding signals when looking at all cryptocurrencies at once. It is challenging to find the reasons causing the herding in such cases since there was no major event that could influence the market [7, 91, 133]. However, this market seems to be mainly driven by investors’ emotions [134] and accessible news on online platforms [21]. In addition, cryptocurrency investors tend to adopt a short-term investment strategy and depend on price movements of the coins rather than reliable factors that are seen in traditional markets [92]. Consequently, the herding in those cryptocurrency communities could arise from either the potential profits or the fear of losing money at that time, encouraging investors to buy or sell. Moreover, investors might be aware of the co-movements between different cryptocurrencies. Thus, when the opportunity to gain profits (or the selling time) arrives, they tend to buy (or sell) the coins that have the same price movements to get higher profits (or mitigate the loss). This could explain why cryptocurrencies mainly formed 2 separate communities in the Pre-Covid-19 (Fig 9) and the third Bull Time (Fig 13) sub-periods, with these communities revealing herding behavior, while there was no herding when we detected the whole cryptocurrency market in the previous section.
@@
Regarding other communities, each of them has a mixture of US ETFs and stocks where the US ETFs tend to act as a basket of most stocks in that community. The herding in these communities follows the same rule as the cryptocurrency communities since each sub-period reveals significant herding behavior in several communities. Remarkably, we obtained several valuable findings from the herding results of these stocks and US ETFs. In particular, there are 4 stock sectors (Energy, Healthcare, Information Technology and Financials) in which for each sector, its assets and relevant ETFs tend to form communities among themselves, throughout the examined period. Moreover, the herding signals found in each of these sectors coincide with the timing of economic and political events that specifically influence that sector.
@@
  • Technology Sector: Technology stocks are grouped into several communities and reveal herding behavior in four sub-periods, including Pre-Covid-19 (coinciding with the US-China trade war [139, 140, 141]), Covid-19 Pandemic (coinciding with the downturn time caused by the pandemic followed by a global economic recession), the second Bull Time (coinciding with the price surge in financial markets[123]) and the second Ukraine-Russia Conflict (coinciding with the AI boom period that gave an impetus to the Technology sector [142, 143]).
@@
  • Healthcare Sector: Herding behavior was found in most healthcare stocks and ETFs during three sub-periods: the Covid-19 Pandemic,the first and third Bull Time. This herding behavior likely stems from the great attraction of investors toward pharmaceutical companies due to their positive performance during the pandemic. In particular, dozens of pharmaceutical companies began vaccine research from early 2020 due to the high demand for vaccinations for the entire world [144, 145, 146]. Moreover, other pharmaceutical companies, though not participating in producing vaccines, produced critical elements that were necessary for vaccine production such as nanoparticles or medical devices for hospitals and daily usage such as the Covid-19 test kits. These contributions also pushed their performance during this time [146, 147].
@@
  • Financial Sector: In this study, the financial sector mainly consists of banking firms. We found that this industry was exposed to herding behavior in the first six sub-periods, except for the second Ukraine-Russia Conflict time. The financial sector is strongly related to economic conditions and reflects the performance of the economy, especially the banking firms since they have the responsibility of preserving the economic stability of a country and also keeping the country away from financial crises through various schemes like reducing interest rates [148], imposing new monetary policy comes from FED [149] and cutting companies’ liquidity risks [150]. Consequently, economic uncertainty forces the financial sector to experience distress. Indeed, the global economy had been experiencing ups and downs continually through out the herding periods in this stock sector.
@@
For the remaining sectors (i.e. Industry, Consumer Discretionary and Consumer Staples), we found that assets within the same sector never group into a large community in all sub-periods, even during the Covid-19 pandemic. Instead, they are broken down and distributed almost equally into several communities. This result indicates that investors did not perform herding on a sector basis in these sectors, even during major events. Furthermore, we acknowledge that some communities are mixed by multiple sectors and reveal herding, but it is unclear why the herding behavior exists in those cases as there seems not to be a rule/convention for these cases. However, most of such communities appear during the economic uncertainty times, i.e. during the trade war (Fig 9) and the Covid-19 pandemic (Fig 10). By contrast, this phenomenon was not pronounced from the first Bull Time sub-period onwards (Figs 11- 15). Thus, it seems to be more likely that negative market conditions resulted in this phenomenon. However, since we don’t have a confident justification for this phenomenon, we will leave this result for future work.
@@
From this experimental result, we can conclude that herding behavior should be observed from two perspectives, one from the entire market and the other from separate subsets (communities) within the market. This is because different perspectives provide us with different aspects of the market. Specifically, herding behavior occurs widely in most assets of a particular type of investment vehicle when there is a critical event that directly impacts the market. Hence, most assets involved are influenced and thus investors who are trading those assets tend to perform herding as a reaction to the event. This is why we found the herding in traditional markets during the US-China trade war, the herding in all markets during the peak of the Covid-19 pandemic and the herding in the cryptocurrency market during bull market conditions, while the remaining sub-periods tend to show no evidence of herding (Tables 3 and 4). By contrast, we found that herding behavior occurs constantly in all types of investment vehicle but at a smaller scale, i.e. community level, as shown in Figs 9-15. In addition, herding behavior found in these communities mainly comes from specific events that solely impact that community. Subsequently, although herding might not exist in a whole market, it can be found by looking at a particular community/sector.
4.3 Contagion Effects between Cryptocurrencies, Stocks and US ETFs
Following the above herding detection results, we conduct another experiment examining the effects of the phenomenon of financial contagion between the assets. This experiment aims to identify the shock drivers between cryptocurrencies, stocks and US ETFs. Since herding has been identified as a cause of the contagion [33], understanding the patterns of contagion over time helps identify the herding transmitters among these assets. Additionally, other authors have also suggested as a link between ETFs and underlying assets, the mechanism of long the stock, short the ETF arbitrage by Hedge funds [151], which tends to amplify the contagion effects between the assets and to give an impetus to herding behavior among investors on not only the ETFs but also their underlying assets.
Based on the 30-min dataset used in this study, we estimate the 30-min lag contagion effects between 222 assets in each of the seven sub-periods using the VAR model. The results are saved in tables and displayed in S1 File. where each column represents a shock receiver and each row represents a shock transmitter, a value that is not 0.0 indicates evidence of contagion from one asset to another asset.
To facilitate the analysis of these tables, we visualize them as 3D figures. In particular, we transform each contagion table result (corresponding to one sub-period) into six figures, including:
  • 1)
    The positive contagion from each of 222 assets to cryptocurrencies.
  • 2)
    The negative contagion from each of 222 assets to cryptocurrencies.
  • 3)
    The positive contagion from each of 222 assets to stocks.
  • 4)
    The negative contagion from each of 222 assets to stocks.
  • 5)
    The positive contagion from each of 222 assets to US ETFs.
  • 6)
    The negative contagion from each of 222 assets to US ETFs.
Examining multiple assets from each type of investment vehicle, rather than focusing only on selected representatives provides a broader view of the contagion effects between different types of investment vehicle. Specifically, the contagion exists widely within each pair of categories, i.e. stocks and US ETFs, cryptocurrencies and US ETFs, and stocks and cryptocurrencies, as several assets in one category transmit shocks to those in another. However, we found that US ETFs are always the strongest shock transmitters to other types of investment vehicle and also among themselves, across different sub-periods. This discloses the strong influence of US ETFs not only on stocks buton also cryptocurrencies, regardless of the existence of economic crises or not. By contrast, although many cryptocurrencies act as shock transmitters to stocks and US ETFs, these contagion effects are generally weak. Without loss of generalization, we display 3D figures of contagion effects between the assets during the Pre-Covid-19 sub-period in Figs 16-21. For each figure, the x-axis represents the assets acting as shock receivers, the y-axis represents the assets acting as shock transmitters, and the z-axis represents the contagion magnitude (absolute value of the contagion coefficient) from one asset to another.
Figure 16: The positive contagion effects from each of 222 assets (y-axis) to cryptocurrencies (x-axis) during the Pre-Covid-19 sub-period. The yellow area represents cryptocurrencies (as shock transmitters) and the orange area represents US ETFs (as shock transmitters). The green spikes represent evidence of positive contagion from one asset to one cryptocurrency.
Figure 17: The negative contagion effects (absolute values) from each of 222 assets (y-axis) to cryptocurrencies (x-axis) during the Pre-Covid-19 sub-period. The yellow area represents cryptocurrencies (as shock transmitters) and the orange area represents US ETFs (as shock transmitters). The red spikes represent evidence of negative contagion from one asset to one cryptocurrency.
Figure 18: The positive contagion effects from each of 222 assets (y-axis) to stocks (x-axis) during the Pre-Covid-19 sub-period. The yellow area represents cryptocurrencies (as shock transmitters) and the orange area represents US ETFs (as shock transmitters). The green spikes represent evidence of positive contagion from one asset to one stock.
Figure 19: The negative contagion effects (absolute values) from each of 222 assets (y-axis) to stocks (x-axis) during the Pre-Covid-19 sub-period. The yellow area represents cryptocurrencies (as shock transmitters) and the orange area represents US ETFs (as shock transmitters). The red spikes represent evidence of negative contagion from one asset to one stock.
Figure 20: The positive contagion effects from each of 222 assets (y-axis) to US ETFs (x-axis) during the Pre-Covid-19 sub-period. The yellow area represents cryptocurrencies (as shock transmitters) and the orange area represents US ETFs (as shock transmitters). The green spikes represent evidence of positive contagion from one asset to one US ETF.
Figure 21: The negative contagion effects (absolute values) from each of 222 assets (y-axis) to US ETFs (x-axis) during the Pre-Covid-19 sub-period. The yellow area represents cryptocurrencies (as shock transmitters) and the orange area represents US ETFs (as shock transmitters). The red spikes represent evidence of negative contagion from one asset to one US ETF.
To explore further the contagion results, we calculate the average contagion magnitude between the three types of investment vehicle considered. In particular, we measure the average contagion magnitude that a type of investment vehicle transmits to other types of investment vehicle and also itself. For example, the value of this metric from stocks to cryptocurrencies is calculated in 2 steps. Firstly, averaging all the absolute contagion coefficients from stocks to each of the cryptocurrencies. After this step, each cryptocurrency has an average contagion magnitude that the stock market transmits to it. Secondly, we average the average values that the stock market transmits to each of the cryptocurrencies.
As shown in Table 5, the contagion from US ETFs to cryptocurrencies is consistently strong and higher than that for other transmissions. Moreover, another strong contagion effect is also seen from US ETFs to stocks. These results remain unchanged across seven sub-periods, indicating a strong influence of US ETFs on both cryptocurrencies and stocks, with a particularly pronounced effect on cryptocurrencies. Similarly, cryptocurrencies are also greatly influenced by the stock market, as indicated by a relatively high contagion magnitude from stocks to cryptocurrencies, although this effect is weaker compared to the influence of US ETFs. In contrast, contagion from cryptocurrencies to traditional assets (both stocks and US ETFs), while present, is negligible across all sub-periods, highlighting the limited influence of cryptocurrencies on these markets. In the case of stocks, they also exhibit only modest contagion effects on US ETFs. Instead, US ETFs primarily transmit shocks among themselves, indicating their strong internal contagion. Notably, contagion effects between these types of investment vehicles peak during the Covid-19 pandemic sub-period, as shown by a rise in most contagion coefficients during this time. After this period, however, these contagion levels declined.
{adjustwidth}
-0.4in0in Transmission Pre-Covid-19 Covid-19 Bull Time 1 Bull Time 2 Bull Time 3 U-R Conflict 1 U-R Conflict 2 Crypto to crypto 0.1674 0.3536 0.2391 0.3040 0.4308 0.1592 0.5855 Stock to stock 0.0962 0.1779 0.1250 0.1201 0.1491 0.1438 0.1377 US ETF to US ETF 0.3191 0.9974 0.4323 0.4951 0.4279 0.4336 0.4974 Stock to crypto 0.4012 0.3823 0.4924 0.8898 0.4939 0.4088 0.3469 US ETF to crypto 1.7424 2.2098 2.3056 3.1864 2.8145 4.7943 1.3000 Crypto to stock 0.0459 0.1225 0.0783 0.0434 0.1264 0.1244 0.0695 US ETF to stock 0.5077 1.0208 0.6153 0.5964 0.7415 0.4669 0.5632 Crypto to US ETF 0.0215 0.0873 0.0293 0.0259 0.1004 0.0811 0.0556 Stock to US ETF 0.0628 0.1562 0.0910 0.1184 0.1091 0.1213 0.1098
Table 5: Average Contagion Magnitude from one type of investment vehicle to the others and itself. Three types of investment vehicle are considered, namely cryptocurrencies, stocks and US ETFs
Linking to the herding behavior discussed in previous sections, US ETFs appear to be one of the herding drivers in financial markets, particularly for cryptocurrencies. In other words, similar trading actions between investors in not only cryptocurrency but also traditional markets may be influenced by the movements of US ETFs, reinforcing the findings of [33]. Additionally, US ETFs seem to exhibit self-reinforcing herding behavior, where they act as drivers of herding among themselves. This substantial amount of contagion (and so influence) from US ETFs to other assets reinforces their role as the central nodes of communities within a correlation network, as shown in the previous section. Moreover, when comparing contagion effects from stocks and cryptocurrencies to US ETFs, stocks generally exert a stronger influence on US ETFs than cryptocurrencies do. To this end, US ETFs and stocks exhibit a greater mutual influence compared to the influence between US ETFs and cryptocurrencies. This explains why US ETFs and stocks are more likely to belong to the same community, while cryptocurrencies tend to form communities between themselves, as discussed in the previous section.
Tables 6 and 7 shows the top 3 US ETFs in each sub-period that transmit their shocks to the highest number of cryptocurrencies and stocks, respectively. As can be seen, the most influential US ETFs tend to be different in different sub-periods, for both cryptocurrencies and stocks. However, SVXY (ProShares Short VIX Short-Term Futures ETF) is the most influential ETF on stocks on a regular basis, transmitting its shocks to the highest number of stocks in the Pre-Covid-19, the second Bull Time and the first Ukraine-Russia Conflict sub-periods. This characteristic is not true for the case of cryptocurrencies since the most influential US ETF varies over time. Furthermore, S&P and Nasdaq-related ETFs tend to be common transmitters to both stocks and cryptocurrencies across different sub-periods.
{adjustwidth}
-0.4in0in Influential Order Pre-Covid-19 Covid-19 Bull Time 1 Bull Time 2 Bull Time 3 U-R Conflict 1 U-R Conflict 2 1 SVXY IYF IDU SVXY QTEC SVXY IWM 2 SLYG QQQ ONEQ FDLO IBB IJT FTXO 3 IJK OEF ITOT QQQ SOXX IDU IYF
Table 6: Top three US ETFs in each sub-period that transmit their shocks to the highest number of stocks
{adjustwidth}
-0.4in0in Influential Order Pre-Covid-19 Covid-19 Bull Time 1 Bull Time 2 Bull Time 3 U-R Conflict 1 U-R Conflict 2 1 QYLD SPY IVE SLYG IJS QQQ FDLO 2 SVXY OEF IYY SVXY FTXO SPY DVY 3 IWB IUSG KBWB IYE KBWB IJK IJH
Table 7: Top three US ETFs in each sub-period that transmit their shocks to the highest number of cryptocurrencies